Fermi National Laboratory
 
Operation Summary for February 20, 2012 - February 22, 2012
 

Monday February 20

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, a Water Group person reported at 9:04 AM that he was making up water for the MI RF Cavity LCW system. A Controls person reported at 10:16 AM that he was rebooting the Switchyard (SY) front end. Water Group personnel reconnected the MI LCW system to CUB at 10:34 AM; the MI LCW system had been in recirculation mode. Operators prepared the SY enclosure G2 and the Neutrino Area enclosure N01 for access at 11:26 AM. Interlock technicians reported at 11:51 AM that they had finished testing G2, N01, NM2, NM3, and NM4. Operations established beam to the Meson MTest area for the EDIT School at 11:52 AM. Operators prepared NM4 for access at noon. Operations halted beam to NuMI at 12:46 PM, opened the MI VCBs, and halted the ramp so that Power Supply personnel could investigate a MI-30 quadrupole magnet power supply. At 1:06 PM, Power Supply personnel reported finding a bad capacitor, several bad resistors, and a problem with the clamp arrangement that connected to the bus. Operators had NM2 and NM3 prepared for access by 3:06 PM. Operations began ramping MI at 3:10 PM; Power Supply personnel reported that they had attached temperature recording labels to the clamps, replaced the traps and contactors in series with the resistors, and brought the power supply back online. Operations resumed sending beam to NuMI at 3:22 PM.

During the evening shift, Operations disabled beam to the MTest EDIT School at 5:47 PM. Operations noticed a problem with the 750 KeV line Buncher at 9:35 PM and called in a Linac person who found the tuning slug for L:IPA2 in local and not responding to the changing tunes of the cavity.

Tuesday February 21

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 12:48 AM, a Linac RF station (LRF3) tripped off and would not come back on; Operators called in a Linac person. Since beam was already off, Operations disabled beam to MiniBooNE at this time in preparation for an access scheduled for the morning. A Linac Klystron RF station (KRF1) tripped off at 1:29 AM; it reset. At 1:39 AM, the Linac person working on LRF3 asked Operations to call in a Linac technician to help him work on the modulator. Linac beam returned around 5:15 AM: Linac personnel reported finding a short from the modulator to the power amplifier caused by a compression ring that had snapped inside the power amplifier. LRF4 tripped off at 6:13 AM; it reset. Operations halted the Linac beam at 6:31 AM to let Linac personnel work on the 750 KeV line Buncher. Beam retuned at 7:23 AM. Operators had MiniBooNE prepared for access by 7:31 AM.

During the day shift, Water Group personnel reported at 8:14 AM that they were making up water to the MI LCW system. Operators reported at 8:17 AM that they had completed a Booster injection tune. Operators had SY enclosure C prepared for access by 8:19 AM. Operations disabled the Linac beam at 8:23 AM so that Linac techs could access the H- Source pit to look for parts to repair the I- Source; beam returned at 8:27 AM. Operators prepare MI-12B for access at 8:33 AM. Operations disabled beam to MI at 8:36 AM so that EE Support personnel could check on the repairs made to the MI-30 quadrupole magnet power supply. Operations had beam back to NuMI by 8:57 AM. Techs in enclosure C came out of access around 10:23 AM. Operators completed another Booster injection tune at 10:40 AM. Linac personnel turned off LRF2 at 11:52 AM for work; Linac resumed sending beam at 12:09 PM. Operations began sending beam to the MTest EDIT School at 12:09 PM. Operations began sending beam to MTA for studies and tuning around 1:05 PM. At 1:41 PM, LRF4 tripped off with no indication of a problem; Linac personnel investigated, replaced an interlock module, and had beam restored about 26 minutes later. Operators prepared the SY G2 enclosure for an access at 1:49 PM.

During the evening shift, Operations established beam to MiniBooNE at 4:42 PM. Operations disabled beam to MTA at 4:59 PM. Operations disabled beam to the EDIT School at 5:42 PM. Operators loaded protons into the Recycler at 10:18 PM. LRF3 tripped off at 10:37 PM; it reset. At 11 PM; LRF2 tripped off; it reset.

Wednesday February 22

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 3:22 AM, BRF11 tripped off; it reset. Operators reported finding four MI ion pumps off (I:IPA604E, I:IPA605A, I:IPA605B, and I:IPA403B) at 4:18 AM; they all reset. The NuMI beam permit tripped off at 4:30 AM due to a MI incomplete RF gap; since no MI RF stations had tripped off, Operators reset the permit. Operators reported at 5:53 AM that they had adjusted the phase set point for LRF5. At 6:14 AM, KRF5 tripped off for the first of two times in 34 minutes; it reset.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo wants the complex to operate with as few beam disruptions as possible for the EDIT School and the other experiments, but there will be a short downtime to check on the MI-30 quadrupole magnet power supply.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/22/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory